# Error mitigation for logical circuits using decoder confidence

### v1

@@ V.3 Numerics
We cannot perform the same simulation for $d=19$ as no logical errors were observed during calibration, but we can estimate its performance with the following model. In the absence of a known distribution of the true log odds, we use the distribution of measured DCS values for $d=19$ and slightly deform it, until the mean $p_{\textnormal{L}}$ associated with this deformed distribution matches the one predicted from our statistics (see Fig. B.1a). We then use Eq. 2 and assume $r$ follows a Gaussian distribution with mean zero and variance $\Delta=1$. The inferred distribution of DCS values closely matches that of the measured ones, as seen in Fig. 8a.
This is therefore an entirely analytic model, in which we have descriptions of the true log odds and the DCS. Moreover, we know the range of log odds associated with each DCS value, so there is no need to calibrate the DCS – in a real scenario, imperfect calibration would indeed be a factor in performance; further work could explore the impact of this. Figure 8d shows impressive improvement in the whole-circuit LEP when we impose an abort condition based on the instantaneous window DCS. Note that the $\Delta=1$ assumption may be pessimistic as an extrapolation from Fig. 4, and with a smaller $\Delta$ the performance is even stronger.
@@
As suggested by the observation in Section IV.2 that most of the logical error risk is contained in a few very risky windows, Fig. 8d shows that the post-selection method does not require exponential overhead for a linear LER decrease (for a significant range of reduction). However, we do observe that the accuracy of the DCS is very significant in the performance of circuit-level exploration of the DCS data. Recall that Section III established a close match between swim distance and the underlying log odds; even so, we see in Fig. 8 panels (c) and (d) that the imperfect nature of the correlation manifests in a considerable reduction in the quality of the information we obtain about the whole-circuit LEP. In panel (d), we show that a modest reduction in the assumed deviation from $\Delta=1$ (somewhat pessimistic given Fig. 4) to $\Delta=0.7$ dramatically improves performance. Conversely, if we were to replace the swim distance with an alternative DCS that is significantly weaker, it is possible that the circuit-level mitigations explored in this paper would become ineffective.
@@ VI Applications
We now discuss applications of the abort protocol and DCS-MLE in quantum phase estimation (QPE) algorithms. In Section VI.1, we perform a detailed resource estimate of our protocols for a Hubbard model ground state energy estimation problem, using statistical QPE. In Section VI.2, we describe a possible use in ancilla-based QPE.
@@ VI.1 Hubbard model resource estimation
In this subsection, we apply our findings to reduce the resource requirements for the implementation of a full logical circuit. Specifically, we find that a $d=21$ surface code can be replaced by $d=19$ to achieve the same LER with a $64\%$ time overhead via the abort protocol. This is a conservative estimate; depending on the metric used to assess logical performance, it could be achieved by MLE without any additional overhead.
@@
We target estimating the ground state energy of the $10\times 10$ Hubbard model to $10^{-2}$ accuracy and at most $10^{-2}$ LER, following the QCELS-based [30] resource estimates in Ref. [31]. The algorithm involves executing $N_{\text{circuits}}=160$ different quantum circuits of various depths, each repeated with $M_{0}=100$ shots to estimate an observable’s expectation value. Appendix E provides more algorithmic details. We estimate that an implementation that does not use the DCS requires a $d=21$ surface code, which corresponds to a LER of approximately $R_{\text{targ}}=8.4\times 10^{-3}$. This is our target LER.
@@
The potential advantage of the DCS arises during the estimation of expectation values. We explore whether the necessary quantum circuits can be executed with $d=19$ (at a starting LER of $R_{\text{start}}=7.7(4)\%$) and still meet the target LER via QEM. Therefore, the desired improvement in the LER is $R_{\text{start}}/R_{\text{targ}}=10.91$. We emulate the behaviour of the $d=19$ code using $d=11$. As discussed in Section IV.2, these distributions can be scaled into each other by adjusting the number $N$ of decoding windows. In Section V.3, we chose the number of decoding windows for the $d=11$ code precisely to approximate the LEP distributions for $d=19$.
@@
By repeatedly sampling $N_{\text{circuits}}M_{0}$ executions from the $\{(\text{LEP},\text{logical error})\}$ distribution at a $60\%$ discard fraction, we estimate that the 95% confidence interval for the effective LER is $0.59\%,R_{\text{targ}}]$, thus very likely achieving the target LER or below, at the cost of $21\%$ total spacetime increase. The latter is calculated in Appendix F, which explains the influence of the code distance on the qubit number and execution duration.
@@
For this algorithm, it may be more useful to analyse the MSPE instead of the effective LER. Using the data shown in Fig. 7, we find that the target MSPE ($0.376\times 10^{-2}$) is achieved by the pure-abort strategy with a similar discard fraction as reported above: $60\%-64\%$. MLE does not give a significant improvement in these cases due to the large Bernoulli variance of $100$ shots. However, in contexts where the bias is more important than the variance, then MLE could achieve the target absolute mean bias ($0.633\times 10^{-2}$) without any time overhead, as shown in Fig. G.1.
@@
To summarise, by exploring the Hubbard model application in a scaled-down form which allows us to apply our earlier empirical results, we believe that in the full problem it should possible to reduce the surface code distance from $21$ to $19$ at a marginal increase in space-time volume. The advantage here would be that the application could run successfully on a smaller quantum computer.
VI.2 Other applications - qualitative remarks
@@
In the case of ancilla-based QPE [32], the final outcome of a noiseless circuit is a bitstring of length $n$ that estimates the first $n$ bits of a desired phase. The noiseless circuit can be run multiple times, resulting in a histogram of output bitstrings; these outputs will get scrambled in the presence of noise. Ref. [33] describes an MLE protocol (based on quasiprobability methods) suitable for deducing the target bitstring. We believe this protocol could be improved using the DCS.
@@
As suggested in Section V.3, we note that DCS-MLE would be very useful in algorithms that do not take full advantage of Heisenberg scaling and instead exploit a trade-off between the number of different circuits run and the number of times each one is repeated [34, 35, 36, 37].
A fundamental question concerning our proposed methods is whether the abort protocol is scalable to large code distances (beyond the early FT era). Most existing works have focused on applications of the DCS with small (<11<11) code distances due to scalability: the sampling overhead of the abort protocol could be unreasonably high. In this work, we provide numerical evidence against this idea by considering the statistical properties of the DCS distribution in Section IV and demonstrating that the method works for a small code with the same LER as a larger one. Ref. [15] brings further evidence towards scalability through a different perspective: under the code capacity noise model, the authors numerically find that, for a fixed abort tolerance, the rate at which the decoder aborts a circuit run exhibits threshold behaviour. Namely, if the physical error rate is below (above) some threshold, the abort rate decreases (increases) as the code distance increases. They find that the threshold is on the order of $10^{-2}$ for code capacity, and expect it to decrease under more realistic noise models. Nevertheless, as long as the abort threshold does not drop below $10^{-4}$, post-selection remains a valuable tool for useful applications in the early fault-tolerant era, such as the quantum phase estimation routine presented here.
@@ VII Discussion and Outlook
Firstly, the tensor network method we used to compute the log success odds was prone to numerical imprecision as there were many samples that led to division by zero or negative odds. For the remaining samples, floating-point precision capped the success log odds values at 16, meaning we were analysing a very restricted subset of cases, especially at large code distances where it is known that the success log odds far exceeds 16.
@@
Thirdly, regarding DCS-based QEM, the calibration curves for large code distances take a long time to obtain because the LEPs are so low. Thus, full calibration curves would have to be provided by the quantum device manufacturers. Nevertheless, Fig. 5b shows that the calibration curves are extremely similar for the same noise model between different code distances, so an approximate calibration curve for $d=19$ could be derived from lower distances. Moreover, if the number of shots is sufficient for MLE, then a calibration curve might not even be necessary, because the extrapolation parameter can also be the raw DCS $\phi$ provided by the decoder; the mitigated expectation value is then the value taken by the fit function in the limit as $\phi\rightarrow\infty$.
Lastly, the linear relation between the noisy outcomes and the noiseless ones applies only to memory experiments. To expand Section V.2 to general logical circuits, we can consider a computational model with layers of logical operations, between which $d$ cycles of stabiliser measurements are performed on each logical qubit. This results in a logical single-qubit Pauli noise model: the logical qubit experiences an $X$, $Y$ or $Z$ flip with a probability given by the DCS. In this model, the corruption function would not be a straight line, but a general polynomial or exponential.
@@ VIII Conclusion
We have shown that the log success odds of a decoding instance can be very well approximated by both the complementary gap and the swim distance for MWPM decoders under phenomenological noise, justifying their status as decoder confidence scores. For a simplified model of window-based decoding, under circuit-level noise, we have shown that DCSs remain rich in information as the circuit size increases, which suggests that they can be used for more complex tasks than state preparation. Finally, we have explored this latter idea in the form of error mitigation techniques for the early fault-tolerant quantum computing era, where aborting and MLE can help run useful applications with a quantum device of limited size. We emphasize that our results rely on the observation that the swim distance is an excellent DCS; the performance of both QEM methods proposed is seen to diminish as the DCS quality decreases. This motivates future work toward even more accurate DCS methods.
@@ Acknowledgements.
We thank Zi-Han Chen for providing code and guidance on their implementation of the swim distance on MWPM. We also thank Christophe Piveteau for providing code for calculating logical coset probabilities via tensor network contraction. For useful discussions, we thank Zhenyu Cai, Armands Strikis and Christopher Chubb. We acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility [42] in carrying out this work and specifically the facilities made available from the EPSRC QCS Hub grant (agreement No. EP/T001062/1). We also acknowledge support from two EPSRC projects: RoaRQ (EP/W032635/1) and SEEQA (EP/Y004655/1). TC and MD acknowledge support from EPSRC DTP studentships.
@@ Appendix A Calculating the Sources of Variation of the DCS
Figure A.1: Decoder confidence score $\phi$ plotted against log success odds $\lambda$ for different code distances $d$. Each blue point is one syndrome sample from the unrotated surface code under a perturbed phenomenological noise model, decoded using MWPM. The best-fit line calculated using Unweighted Least Squares is shown in red; the height of the shaded region is the root mean squared residual, and is plotted as $\sigma_{r}$ in Fig. 4. Due to floating-point precision the success log odds values are capped at 16, so we ignore samples whose complementary gap exceeds 16.
@@
$\widehat{\sigma_{\alpha(\lambda)}}=\sqrt{s_{\phi}^{2}-s_{r}^{2}}$ (8)
@@ Appendix B LEP Variance of a Many-Window Experiment
$\mu_{N,d}$ $=\tfrac{1}{2}[1-(1-2\mu_{1,d})^{N}],$ (9)
$\sigma_{N,d}^{2}$ $=\tfrac{1}{4}\left\{[(1-2\mu_{1,d})^{2}+(2\sigma_{1,d})^{2}]^{N}-(1-2\mu_{1,d})^{2N}\right\}.$ (10)
@@ C.1 Abort protocol
Here, we describe how we generate datasets for the LEP and logical error of each an $N$-window circuit. We perform simulations up to a code distance of $11$, which is the largest distance at which enough logical errors could be detected during calibration within $10^{8}$ shots. To avoid simulating deep circuits and calculating their DCSs, we emulate the behaviour of a circuit comprising $N=$2.38\text{\times}{10}^{5}$\$ windows by sampling from a pool of single-window circuits, each characterised by a calibrated LEP $p_{i}$ and a logical error $x_{i}\in\{0,1\}$, where $x_{i}=0$ indicates successful decoding. The $N$-window circuit is considered to have experienced an error if the decoding of an odd number of the $N$ windows failed: $X=1$ if $\sum_{i=1}^{N}x_{i}$ is odd, else $X=0$. The assigned LEP for an $N$-window circuit is given by Eq. 3.
Next, we explain how the abort protocol affects the LEP distribution. Let the single-window abort rate be $\rho$, achieved by setting some LEP threshold. This means that, in an $N$-window circuit, each window has an LEP larger than that threshold with probability $\rho$. The probability to discard the circuit by aborting at any point throughout its run is $f=1-(1-\rho)^{N}$. Throughout this work, we refer to $f$ as the discard fraction.

### v2

@@ V.3.2 Distance-19 Surface Code
We cannot perform the same simulation for $d=19$ as no logical errors were observed during calibration, but we can estimate its performance with the following model. In the absence of a known distribution of the log success odds $\lambda$, we use the distribution of measured DCS values for $d=19$ and slightly deform it, until the mean $p_{\textnormal{L}}$ associated with this deformed distribution matches the one predicted from our statistics (see Fig. B.1a). We then use Eq. 2 and assume $r$ follows a Gaussian distribution with mean zero and variance $\Delta=1$. The inferred distribution of DCS values closely matches that of the measured ones, as seen in Fig. 8a. This is therefore an entirely analytic model, in which we have descriptions of $\lambda$ and the DCS. Moreover, we know the range of $\lambda$ associated with each DCS value, so there is no need to calibrate the DCS – in a real scenario, imperfect calibration would indeed be a factor in performance; further work could explore the impact of this.
@@
Figure 8d shows impressive improvement in the whole-circuit LEP when we impose an abort condition based on the instantaneous window DCS. As suggested by the observation in Section IV.2 that most of the logical error risk is contained in a few very risky windows, we see the post-selection method does not require exponential overhead for a linear decrease in LER (for a significant range of reduction). However, we do observe that DCS accuracy is very significant in the amount of reduction. Recall that Section III established a close match between the swim distance and $\lambda$; even so, we see in Fig. 8c,d that the imperfect correlation manifests in a considerable worsening of the quality of the information we obtain about the whole-circuit LEP. Note that the $\Delta=1$ assumption is somewhat pessimistic as an extrapolation from Fig. 4, and with a modest reduction to $\Delta=0.7$ the performance dramatically improves. Conversely, if we were to replace the swim distance with an alternative DCS that is significantly weaker, it is possible that the whole-circuit LEP mitigations explored in this paper would become ineffective.
@@ VI Applications
We now discuss applications of the abort protocol and DCS-MLE in quantum phase estimation (QPE) algorithms. In Section VI.1, we perform a detailed resource estimate of our protocols for a Hubbard model ground state energy estimation problem, using statistical QPE. In Section VI.2, we qualitatively describe possible uses of DCS-MLE in other applications.
@@ VI.1 Hubbard Model Resource Estimation
In this subsection, we apply our findings to reduce the resource requirements for the implementation of a full logical circuit. Specifically, we find that the surface code distance can be reduced from 21 to 19 and achieve the same LER with a $64\%$ time overhead via the abort protocol. This is a conservative estimate; depending on the metric used to assess logical performance, it could be achieved by MLE without any additional overhead.
@@
We target estimating the ground state energy of the $10\times 10$ Hubbard model to $10^{-2}$ statistical accuracy (in absence of physical noise) and at most $10^{-2}$ LER (from noise added on top of statistical fluctuations), following the QCELS-based [30] resource estimates in Ref. [31]. The algorithm involves executing $N_{\text{circuits}}=160$ different quantum circuits of various depths, each repeated with $M_{0}=100$ shots to estimate an observable’s expectation value. This means that $N_{\text{executions}}=N_{\text{circuits}}M_{0}$ circuit executions would take place in practice. Appendix E provides more algorithmic details. We estimate that an implementation that does not use the DCS requires a $d_{\text{targ}}=21$ surface code, which corresponds to a LER of approximately $R_{\text{targ}}=8.4\times 10^{-3}$; this is our target LER.
@@
The potential advantage of the DCS arises during the estimation of expectation values. We explore whether the necessary quantum circuits can be executed with a smaller $d_{\text{small}}=19$ code (at an unmitigated LER of $R_{\text{small}}=7.7(4)\%$) and still meet the target LER via QEM. Therefore, the desired improvement in the LER is $R_{\text{small}}/R_{\text{targ}}=10.91$. We emulate the behaviour of the $d=d_{\text{small}}$ code using $d=11$ and 1. In Section V.3.1, the number of decoding windows was chosen precisely to approximate the LEP distributions for $d=d_{\text{small}}$.
@@
By repeatedly sampling $N_{\text{executions}}$ times from the emulated experiment distribution used in Section V.3.1 at a $60\%$ discard fraction, we estimate that the 95% confidence interval for the effective LER is $0.59\%,R_{\text{targ}}]$, thus achieving the target LER or better, at a $21\%$ increase in spacetime volume (calculated in Appendix F).
@@
For this algorithm, it may be more useful to analyse the MSPE instead of the effective LER. Using the data shown in Fig. 7, we find that the target MSPE ($0.376\times 10^{-2}$) is achieved by the pure-abort strategy with a similar discard fraction as reported above: $60\%$. MLE does not give a significant improvement in these cases due to the large Bernoulli variance of $100$ shots. However, in contexts where the bias is more important than the variance, MLE could achieve the target without any time overhead, as discussed in Appendix G.
@@
To summarise, we explore the Hubbard model application in a scaled-down form which allows us to apply our earlier empirical results from Section V.3.1. We believe that for the full problem it should possible to reduce the surface code distance from $21$ to $19$ at a marginal increase in space-time volume. The advantage here is that the application could run successfully on a smaller quantum computer.
VI.2 Other Applications
@@
In ancilla-based QPE [32], the final outcome of the noiseless circuit is a bitstring of length $n$ that estimates the first $n$ bits of a desired phase. The noiseless circuit can be run multiple times, resulting in a histogram of output bitstrings; these outputs will get scrambled in the presence of noise. Ref. [33] describes an MLE protocol (based on quasiprobability methods) suitable for deducing the target bitstring. We believe this protocol could be improved using the DCS.
@@
As suggested in Section V.3.1, DCS-MLE would be very useful in algorithms that do not take full advantage of Heisenberg scaling and instead exploit a trade-off between the number of different circuits run and the number of times each one is repeated [34, 35, 36, 37].
@@ VII Discussion and Outlook
Firstly, the tensor network method we used to compute $\lambda$ was prone to numerical imprecision as there were many samples that led to division by zero or negative odds. For the remaining samples, floating-point precision capped the $\lambda$ values at 16, meaning we were analysing a very restricted subset of cases, especially at large code distances where it is known that $\lambda$ can far exceed 16.
@@
Thirdly, regarding DCS-based QEM, the calibration curves for large code distances take a long time to obtain because the LEPs are so low. Thus, full calibration curves would have to be provided by the quantum device manufacturers. Nevertheless, Fig. 5b shows that the calibration curves are extremely similar for the same noise model between different code distances, so an approximate calibration curve for $d=19$ could be derived from lower distances.
Lastly, we emulated deep logical circuits by stitching together the results from individual memory experiments. A more comprehensive simulation would account for multiple interacting logical qubits, meaning the corruption function $q$ would in general be nonlinear in $P_{\textnormal{L},j}$. Note that a single-logical-qubit Pauli noise model can be constructed for general circuits based on the DCS. We can consider a computational model with layers of logical operations, between which $d$ cycles of stabiliser measurements are performed on each logical qubit. Hence, each logical qubit experiences an $X$, $Y$, or $Z$ error with a probability given by the DCS.
A fundamental question concerning our proposed methods is whether the abort protocol is scalable to large code distances (beyond the early FT era). Most existing works have focused on applications of the DCS with small (<11<11) distances due to scalability: the sampling overhead of the abort protocol could be unreasonably high. Section IV provides numerical evidence against this idea by considering the statistical properties of the DCS distribution and demonstrating that the method works for a small code with the same LER as a larger one. Ref. [15] provides alternative evidence towards scalability: the authors find numerically that the rate at which the decoder aborts a circuit run exhibits threshold behaviour. Namely, for a fixed abort tolerance, if the physical error rate is below (above) some threshold, the abort rate decreases (increases) as the code distance increases. They observe that the threshold is on the order of $10^{-2}$ for the code capacity noise model, and expect it to decrease under more realistic noise models. Nevertheless, as long as the abort threshold does not drop below $10^{-4}$, post-selection remains a valuable tool for useful applications in the early fault-tolerant era, such as the QPE routine described in Section VI.2.
@@ VIII Conclusion
We have shown that the log success odds of a decoding instance can be very well approximated by both the complementary gap and the swim distance for MWPM under phenomenological noise, justifying their status as DCSs. For a simplified model of window-based decoding under circuit-level noise, we have shown that DCSs remain rich in information as the circuit size increases, which suggests that they can be used for more complex tasks than state preparation. Finally, we have explored this latter idea in the form of QEM techniques for the early fault-tolerant quantum computing era, where aborting and MLE can help run useful applications with a quantum device of limited size. We emphasise that our results rely on the observation that the swim distance is an excellent DCS; the performance of both proposed QEM methods diminishes as the DCS quality decreases. This motivates future work toward even more accurate DCS methods.
@@ Acknowledgements.
We thank Zi-Han Chen for providing code and guidance on their implementation of the swim distance on MWPM. We also thank Christophe Piveteau for providing code for calculating logical coset probabilities via tensor network contraction. For useful discussions, we thank Zhenyu Cai, Armands Strikis and Christopher Chubb. We acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility [42] in carrying out this work and specifically the facilities made available from the EPSRC QCS Hub grant (agreement No. EP/T001062/1). We also acknowledge support from two EPSRC projects: RoaRQ (EP/W032635/1) and SEEQA (EP/Y004655/1). TC and MD acknowledge support from EPSRC DTP studentships. TC was supported by JST ASPIRE Japan Grant Number JPMJAP2319.
Author Contributions
TC conceived and performed the tensor network comparison in Section III. MD designed the DCS-MLE procedure and performed the analyses in Sections IV and V.3.1, and the resource estimation in Section VI.1. SCB performed the analysis in Section V.3.2 and supervised the project. All authors contributed equally to writing the manuscript.
@@
Shortly after the first version of this manuscript was announced, another related preprint [44] was announced that proposes a logical error mitigation protocol called SALEM, conditioned directly on the decoder syndrome. The authors treat different syndromes as inducing different logical error channels, and syndrome-dependent mitigation or rejection strategies are applied to reduce logical errors at the channel level. While both our work and theirs exploit run-to-run variability in LEP and abort risky circuit executions based on the DCS, they differ substantially in methodology. For accepted executions, SALEM mitigates errors via inverting logical error channels, whereas our approach further uses the DCS as a parameter in classical post-processing via maximum-likelihood estimation. The requirements to implement the methods are also different: SALEM requires channel characterisation, while DCS-MLE requires a calibrated mapping between the window-level DCS and the LEP.
@@ Appendix A Calculating the Sources of Variation of the DCS
Figure A.1: Decoder confidence score $\phi$ plotted against log success odds $\lambda$ for different code distances $d$. Each blue point is one syndrome sample from the unrotated surface code under a perturbed phenomenological noise model, decoded using MWPM. The best-fit line calculated using Unweighted Least Squares is shown in red; the height of the shaded region is the root mean squared residual, and is plotted as $\sigma_{r}$ in Fig. 4. Due to floating-point precision the $\lambda$ values are capped at 16, so we ignore samples whose complementary gap exceeds 16.
@@
$\widehat{\sigma_{\alpha(\lambda)}}=\sqrt{s_{\phi}^{2}-s_{r}^{2}},$ (9)
@@ Appendix B LEP Variance of a Many-Window Experiment
$\mu_{N,d}$ $=\tfrac{1}{2}\big[{1-(1-2\mu_{1,d})^{N}}\big],$ (10)
$\sigma_{N,d}^{2}$ $=\tfrac{1}{4}\big\{{\big[{(1-2\mu_{1,d})^{2}+(2\sigma_{1,d})^{2}}\big]^{N}-(1-2\mu_{1,d})^{2N}}\big\}.$ (11)
@@ C.1 Abort Protocol
Here, we describe how we generate datasets for the LEP and logical error of each $N$-window circuit. We perform simulations up to a code distance of $11$, which is the largest distance at which enough logical errors could be detected during calibration within $10^{8}$ shots. To avoid simulating deep circuits and calculating their DCS values, we emulate the behaviour of a circuit comprising $N=$2.38\text{\times}{10}^{5}$\$ windows by sampling from a pool of single-window circuits, each characterised by a calibrated LEP $p_{i}$ and a logical error $x_{i}\in\{0,1\}$, where $x_{i}=0$ indicates successful decoding. The $N$-window circuit is considered to have experienced an error if the decoding of an odd number of the $N$ windows failed: $X=1$ if $\sum_{i=1}^{N}x_{i}$ is odd, else $X=0$. The assigned LEP for an $N$-window circuit is given by Eq. 3.
Next, we explain how the abort protocol affects the LEP distribution. Let the single-window abort rate be $\rho$, achieved by setting some LEP threshold. This means that, in an $N$-window circuit, each window has an LEP larger than that threshold with probability $\rho$. The probability to discard the circuit by aborting at any point throughout its run is $f=1-(1-\rho)^{N}$. Throughout this work, we refer to $f$ as the discard fraction. This is analogous to Fig. 8d.